Solve the following Initial Value Problem using (i)R-K method of O(h2)

and (ii) R-K method of O(h4)

y' = x2y + x3 and y(0) = 1.

Find y(0.4) taking h = 0.2, where y' means dy/dx




1
Expert's answer
2021-10-29T03:00:21-0400

(i) Rung-Kutta method of second orde

"y'=x^2y+x^3=\frac{dy}{dx}"

"h=0.2"

"y(0)=1, y_0=1, x_0=0"

"f(x,y)=x^2y+x^3, u_1=u_0+h=0+0.2\implies u_1=0.2"

Applying Runge-Kutta order 2,

"k_1=hf(x_0,y_0)"

"=0.2f(0,1)=0.2(0^2+0^3)=0"

"k_2=hf(x_0+h,y_0+k_1)=hf(0.2,1)"

"=0.2(0.2^2\cdot1+0.2^3)=0.0096"

"\therefore y_1=y(0.2)=y_0+\frac{1}{2}(k_1+k_2)"

"=1+\frac{1}{2}(0+0.0096)=1.0048"

Now; "x_1=0.2, y_1=1.0048,h=0.2"

"\therefore u_2=u_1+h=0.2+0.2=0.4"

"f(x,y)=x^2y+x^3"

Applying R-K method order 2

"k_1=hf(x_1,y_1)=0.2f(0.2,1.0048)"

"=0.2((0.2)^2\cdot1.0048+(0.2)^3)=0.0096"

"k_2=hf(x_1+h,y_1+k)"

"=hf(0.4,1.0144)"

"=(0.2)((0.4)^2\cdot1.0144+(1.0144)^3)=0.2412"

"y_2=y(0.4)=y_1+\frac{1}{2}(k_1+k_2)"

"=1.0048+\frac{1}{2}(0.0096+0.2412)"

"=1.1302"

(ii) "f(x,y)=x^2y+x^3"

"h=0.2"

"y(0)=1, y_0=1, x_0=0"

Applying Runge-Kutta method of fourth order,

"k_1=hf(x_0,y_0)=0.2f(0,1)"

"=0.2(0+0)=0"

"k_2=hf(x_0+\frac{h}{2},y_0+\frac{k}{2})=hf(0.1,1)"

"=0.2(0.1^2\cdot1+(0.1)^3)=0.0022"

"k_3=hf(x_0+\frac{h}{2},y_0+\frac{k_1}{2})=hf(0.1,1.0011)"

"=0.2((0.1)^2\cdot(1.0011)+(0.1)^3)=0.0022"

"k_4=hf(x_0+h,y_0+k_3)=hf(0.2,1.0022)"

"=(0.2)((0.2)^2\cdot(1.0022)+(0.2)^3)=0.0096"

"k=\frac{1}{6}(k_1+2k_2+2k_3+k_4)=0.0031"

"\therefore y_1=y(0.2)=y_0+k=1+0.0031=1.0031"

Now "x_1=0.2,y_1=1.0031,h=0.2"

"f(x,y)=x^2y+x^3"

"x_2=x_1+h=0.2+0.2=0.4"

By applying R-K method of order 4,

"k_1=hf(x_1,y_1)=0.2f(0.2,1.0031)"

"=(0.2)((0.2)^2\cdot(1.0031)+(0.2)^3)"

=0.0096

"k_2=hf(x_1+\frac{h}{2},y_1+\frac{k_1}{2})=hf(0.3,1.0079)"

"=(0.2)((0.3)^2\cdot(1.0079)+(0.3)^3)=0.0235"

"k_3=hf(x_1+\frac{h}{2},y_1+\frac{k_2}{2})=hf(0.3,1.0148)"

"=(0.2)((0.3)^2\cdot(1.0148)+(0.3)^3)=0.0237"

"k_4=hf(x_1+h,y_1+k_3)=hf(0.4,1.0268)"

"=(0.2)((0.4)^2\cdot(1.0268)+(0.4)^3)=0.0456"

"k=\frac{1}{6}(k_1+2k_2+2k_3+k_4)"

"=0.0249"

"\therefore y_2=y(0.4)=y_1+k"

"=1.0031+0.0249"

"=1.028"